Equipment acquired at a cost of $70,000 has an estimated residual value of $4,000 and an...

  1. Equipment acquired at a cost of $70,000 has an estimated residual value of $4,000 and an estimated useful life of 10 years. It was placed into service on October 1 of the current fiscal year, which ends on December 31.

    If necessary, round your answers to the nearest cent.

    a. Determine the depreciation for the current fiscal year and for the following fiscal year by the straight-line method.

    Depreciation
    Year 1 $
    Year 2 $

    b. Determine the depreciation for the current fiscal year and the following fiscal year by the double-declining-balance method.

    Depreciation
    Year 1 $
    Year 2 $
Answer #1

Answers

  • Workings

A

Cost

$            70,000.00

B

Residual Value

$              4,000.00

C=A - B

Depreciable base

$            66,000.00

D

Life [in years]

10

E=C/D

Annual SLM depreciation

$              6,600.00

F=E/C

SLM Rate

10.00%

G=F x 2

DDB Rate

20.00%

  • Requirement [a]

Depreciation

Year 1

$ 1650 [$6600 x 3months/12]

Year 2

$ 6600

  • Requirement [b]

Depreciation

Year 1

$ 3500

Year 2

$ 13300

--Working

Year

Beginning Book Value

Depreciation rate

Depreciation expense

Ending Book Value

1

$              70,000.00

20.00%

$            3,500.00 [70000 x 20% x 3/12]

$             66,500.00

2

$              66,500.00

20.00%

$         13,300.00

$             53,200.00